Mr. Takahashi has n = 16 and he is trying to compute a 90% confidence interval. However, he does not have access to the population’s standard deviation and instead he is using a sample standard deviation. Mr. Takahashi should use a _-stat and it should be ___.

A

Z; 1.753

B

t; 1.753

C

t; 1.746

D

Z; 1.746
Answer #1

Solution :

Given that,

sample size = n = 16

Degrees of freedom = df = n - 1 = 16 - 1 = 15

At 90% confidence level

= 1 - 90%

=1 - 0.90 =0.10

/2 = 0.05

t/2,df = t0.05,15 = 1.753

t = 1.753

correct option is = B